One person swings on a swing and finds that the period? T0 is equal to 3.0s A second person of equal mass joins him on the same wing. With two people swinging, what is the new period of oscillation, T(new) ?

A- T(new) = 6.0s
B- 3.0s < T(new) < 6.0s
C- T(new) = 3.0s
D- 1.5s < T(new) < 3.0s
E- there is not sufficient information to determine T(new)
Answer #1

To = 3 sec

Tnew = ?

Io = moment of inertia of one person

Inew = moment of inertia of two person = 2 Io

using conservation of angular momentum

Io Wo = Inew Wnew

Io (2\pi/To) = Inew (2\pi/Tnew)

Io (2\pi/To) = (2Io) (2\pi/Tnew)

1/To = 2 /Tnew

Tnew = 2 To = 2 x 3 = 6 sec